Use the following functions that show quantity demanded and quantity supplied to answer the question.

Qd = 600 - 30p

Qs = -300 + 120p


If the price is R8, then there will be a surplus of _____ units. 



1
Expert's answer
2021-06-07T10:36:05-0400

"Qs = Qd"

"600-30P = -300 +120p"

"600+300 = 120p + 30p"

"900 = 150p"

"P = 900\/150"

"= 6"

"R8 = 6 \u00d7 8"

"= 48 units"